OpenStudy (anonymous):

can someone help me simplify this: the quantity 10 times x to the 6th power times y to the third power plus 20 times x to the third power times y to the 2nd power all over 5 times x to the third power times y

OpenStudy (shane_b):

\[\frac{10x^6y^3+20x^3y^2}{5x^3y}\] Simplify this?

OpenStudy (anonymous):

yes please

OpenStudy (shane_b):

First, break it into two fractions:\[\frac{10x^6y^3}{5x^3y}+\frac{20x^3y^2}{5x^3y}\] To simplify each fraction, use the exponent rule which states that when you divide exponents with the same base, you subtract the exponent in the denominator from the exponent in the numerator. If you do this, what do you get?

OpenStudy (shane_b):

I'll do the first one for you to make sure it's clear. BTW: You also need to divide the coefficients normally.\[\frac{10x^6y^3}{5x^3y}=2x^3y^2\]

OpenStudy (anonymous):

dude i just need the answer im being timed

OpenStudy (shane_b):

I'm sure someone else here will help you then. Good luck.

OpenStudy (anonymous):

thanks for nothing

OpenStudy (anonymous):

2x^3y^2 + 4y^2, got that?